OpenStudy (anonymous):

find the ratio in which the point (7/9,-2/9) divides the interval joining (3,-6) and (-2,7)

OpenStudy (amistre64):

given 2 point (a,b) and (p,q); any point dividing (x,y) creates distances between them such that: \[\sqrt{(a-x)^2+(b-y)^2}+\sqrt{(p-x)^2+(q-y)^2}=\sqrt{(a-p)^2+(b-q)^2}\]